OpenStudy (foxythepirate0984):

students in classes, displayed below, ate the same ratio of cheese pizza slices to pepperoni pizza slices. Below is the Slices of Cheese Pizza column from the table. Fill it in with fractions or mixed numbers (like 7 1/2) Slices of Cheese Pizza

OpenStudy (kittiwitti1):

Someone likes FNaF.

OpenStudy (kittiwitti1):

Haha.... where is this table you're referring to?

OpenStudy (foxythepirate0984):

im getting it now

OpenStudy (kittiwitti1):

Alright

OpenStudy (kittiwitti1):

This is an addition thing right?

OpenStudy (foxythepirate0984):

uh. i dont know , if your asking like fraction then its fractions

OpenStudy (kittiwitti1):

Like this? http://icecream.me/5a4556752d67d607d8cc49f8da8ef5c6

OpenStudy (foxythepirate0984):

yes

OpenStudy (kittiwitti1):

Well okay then, you can do this yourself right? :p

OpenStudy (foxythepirate0984):

not really. it seems hard and i missed the classes. :(

OpenStudy (kittiwitti1):

:T Ok here's a hint: http://icecream.me/354b6278c7f17e5d77bbbf35cbdefe61

OpenStudy (kittiwitti1):

Whoops don't write 30 in that box, I meant 21

OpenStudy (foxythepirate0984):

i still dont get what to do

OpenStudy (kittiwitti1):

Just stick random fractions in the Pepperoni boxes first

OpenStudy (foxythepirate0984):

k

OpenStudy (kittiwitti1):

Wait... Okay here's a tip. Fill in the first row: Pepperoni 3/4, Cheese 6 and 1/4. What does that add up to?

OpenStudy (foxythepirate0984):

your saying something wrong cause in the picture you sent me, it fills out almost all the colums, when i need to fill out the cheese pizza i think

OpenStudy (foxythepirate0984):

6 5/4

OpenStudy (kittiwitti1):

You need to fill out the columns to get the cheese pizza lol. No, it's 7\[\frac{3}{4}+6\frac{1}{4}=6+\frac{1+3}{4}=6\frac{4}{4}=7\]

OpenStudy (kittiwitti1):

Sorry I messed up on the fractions. But yeah it's 7.

OpenStudy (foxythepirate0984):

mk, so the answer is 7?

OpenStudy (kittiwitti1):

That's the answer for total pizza. So basically we keep going like that per row, until you've filled all the columns like that then just take the cheese pizza column and BAM there's your answer! The first row has "total pizza = 7", I just put two fractions that added up to 7 in the pepperoni and cheese sections

OpenStudy (foxythepirate0984):

i dont understand

OpenStudy (kittiwitti1):

First row: Cheese and Pepperoni slices are blank. You have a total pizza slice count of 7. The fractions I gave you add up to 7, so that fills the first row.

OpenStudy (foxythepirate0984):

im still confused :(

OpenStudy (kittiwitti1):

http://prntscr.com/8rhko8 x + y = z. In row one, we have z=7 x and y must be fractions or mixed numbers.

OpenStudy (kittiwitti1):

I made x=3/4, that's a fraction and I made y = 6 and 1/4, that's a mixed number. Together, those add up to 7. x+y=z, yes?